OpenStudy (nathaliama13):

A zoo train ride costs $3 per adult and $1 per child. On a certain day, the total number of adults (a) and children (c) who took the ride was 30, and the total money collected was $50. What was the number of children and the number of adults who took the train ride that day, and which pair of equations can be solved to find the numbers?

OpenStudy (nathaliama13):

Options: A. 20 children and 10 adults Equation 1: a + c = 30 Equation 2: 3a + c = 50 B. 10 children and 20 adults Equation 1: a + c = 30 Equation 2: 3a − c = 50 C. 20 children and 10 adults Equation 1: a + c = 30 Equation 2: 3a − c = 50 D. 10 children and 20 adults Equation 1: a + c = 30 Equation 2: 3a + c = 50

OpenStudy (anonymous):

you use the solving equations method , u could do substitution or elimination.

OpenStudy (nathaliama13):

A or D? @Flawless01

OpenStudy (nathaliama13):

Or A or C?

OpenStudy (anonymous):

did u get the answer? @nathaliaMA13

OpenStudy (nathaliama13):

That was 8th grade homework...I'm a Freshman now. I think I solved it but I don't remember what I got. @rolotherabbit1